Please help me with the following problem:
The question is as followed:
Explain the steps that you need to take in order to resolve the system of equations as followed in applying the method of elimination. You don't have to resolve the problem itself.
1- 5x + 2y =5
2- 3x + 4y= -23
Thank you

Answer by checkley71(8403)

5X+2Y=5
3X+4Y=-23
IF YOU MULTIPLY THE FIRST EQUATION BY -2 WE GET
-10X-4Y=-10 NO ADD TO THE SECOND EQUATION
3X+4Y=-23
---------------------
-7X=33
X=-33/-7
X=33/7
5*33/7+2Y=5
165/7+2Y=5
2Y=5-165/7
2Y=(35-165)/7
2Y=-130/7
Y=(-130/7)/2
Y=-130/7*1/2
Y=-130/14
PROOF
5*33/7+2*-130/14=5
165/7-260/14=5
(330-260)/14=5
70/14=5
5=5